# Semialgebraic Neural Networks: From roots to representations

### v1

@@ E.3 Proof of Theorem LABEL:thm-continuous-semialgebraic-functions-computable-by-homotopy
Let $F:\mathbb{R}^{m}\to U_{F}$ be a given continuous semialgebraic function. For every ${x\in\mathbb{R}^{m}}$ there exists $H_{x,a}$ with the form equation 11 such that for almost every ${a\in U_{F}}$, the following hold:
@@
We now consider each hypothesis of that theorem in turn:
@@ E.4 Proof of Theorem 12
All that remains it to prove this procedure can be repeated until time $s=1$ is reached. Due to the clamping in in line 5 of algorithm 1, $z$ remains in $[-c_{\max},c_{\max}]^{n+1}$ for time $s\in[0,1]$. Since $\mathcal{N}\in ISDnet(m,n,1)$ is a piecewise polynomial, it is Lipschitz on the compact domain $\{x\}\times[-c_{\max},c_{\max}]$, which means there exists $\epsilon>0$ such that for each step $\tau$ of the $s$-glueing procedure above, we can choose $s_{\tau+1}$ such that $s_{\tau}+\epsilon<s_{\tau+1}$, and the conclusion follows. ∎
@@ E.5 Discontinuous functions
Furthermore, we carefully construct each piece of the computation to lie in disjoint $(t,s)$-space, so they can be glued together in a way consistent with a single continuous ISD network $\mathcal{N}$. This $\mathcal{N}$ computes $F$ while solving the ODE (7)-(8). Ensuring that the required $\mathcal{N}$ is indeed an ISD network (rather than an arbitrary continuous one) is one of the primary challenges in these constructions.
@@ E.6 Characteristic functions
Hypothesis 67 guarantees the sets
@@
$\texttt{ODESolve}(\mathcal{N}^{\prime}_{2},c,x,(0,1,\frac{1}{2}),1)=(0,1)-(0,\chi_{S_{2}}(x)).$ (81)
@@
$\texttt{ODESolve}(\mathcal{N}^{\prime}_{2},c,x,(0,0,\frac{1}{2}),1)=(0,0),$ (82)
@@
We can now handle the case of an arbitrary semialgebraic set $S$. The key is to exploit the “cylindrical decomposition” of $S$: every semialgebraic set in $\mathbb{R}^{m}$ is the union of finitely many disjoint sets homeomorphic to $(-1,1)^{d}$, $d\leq m$ Bochnak et al. [1998]. In particular, it is the union of finitely many disjoint locally closed sets. Since locally closed sets are the intersection of an open and a closed set, we have already shown how to represent their characteristic functions with SANNs. Furthermore, since these locally compact sets are disjoint, the characteristic function of their union can be represented by summing the individual characteristic functions with $s$-glueing. ∎
@@ E.7 Proof of Theorem 13
Every semialgebraic function has finitely many connected components [Bochnak et al., 1998], so let $\{S_{j}\}_{j=1}^{K}$ be a semialgebraic decomposition of $\mathbb{R}^{m}$ such that $F$ is continuous on each $S_{i}$.
@@
Using 42, modify each $\mathcal{N}_{i}$, $c_{i}$, so that integration occurs in the interval $s\in[1/2,1]$. Furthermore, using 43, modify each $\mathcal{N}_{i}$, $c_{i}$ to begin integration at $(0,3j).$ Finally, $t$-glue these modified $\mathcal{N}_{j}$, and $s$-glue the result to $\mathcal{N}_{0}$ at $s=1/2$. ∎
@@ E.9 SANNs always compute semialgebraic functions
In our considerations, we use the Tarski–Seidenberg theorem has several equivalent formulations. We use the following:
@@ Appendix F Numerical example: Inverse problem for electrical resistor networks
Suppose we are given an electrical network $\Gamma=(V,E)$ with vertices $V$ and edges $E\subset V\times V.$ Physically, an electrical network is a system of connected resistors and it is modeled by a graph where all edges are resistor. An inverse problem for an electrical network is to find one or all graphs and resistors on the edges when the outcome of all possible voltage and current measurements in some of vertices of the graph (called the boundary nodes) are given, see Figure 5.
We partition $V$ into boundary nodes $V_{B}$ and interior nodes $V_{I}$. Further let $\gamma:E\to\mathbb{R}_{\geq 0}$ be the edge conductivities. The network response matrix (i.e. Dirichlet-to-Neumann map) $\Lambda\in\mathbb{R}^{|V_{B}|\times|V_{B}|}$ computes the current at the boundary nodes resulting from applying voltages on $V_{B}$. Our task is to recover the conductivites $\gamma$ from measurements $\Lambda$ of voltage-response on the boundary. See Curtis and Morrow [2000] for a thorough treatment of this inverse problem, and Forcey [2023] for recent developments.
@@
$K$ is a symmetric matrix with conductivities on off-diagonals and that satisfies Kirchhoff’s Law—the sum of entries in each row is 00. Curtis and Morrow [2000] showed that $\Lambda$ is the Schur complement of a particular submatrix of the Kirchhoff matrix. Precisely, we write
@@
The inverse problem is to recover an element of the preimage $F^{-1}(\{\Lambda\})$. Since preimages of semialgebraic sets (in this case the single point $\{\Lambda\}$) of semialgebraic functions ($F$) are themselves semialgebraic [Bochnak et al., 1998], we conclude $F^{-1}(\{\Lambda\})$ is a semialgebraic set.
In certain special network topologies, such as for rectangular (Figure 5), the preimage of $F^{-1}(\Lambda)$ is a singleton, that is, the values $\gamma(i,j),$ $(i,j)\in E$ of the resistors on the graph $(V,E)$ are uniquely determined by the voltage-to-current map $\Lambda$, see [Curtis and Morrow, 2000].
@@
This inverse problem has had a wide impact in several fields. In medical imaging it has led to new imaging modalities, for example, in monitoring lung function. Recently, EIT is used in monitoring the lungs of the COVID-19 patients, see Cappellini et al. [2024].
@@
We demonstrate the feasibility of training SANNs to perform non-trivial tasks using the network EIT problem as an example. Specifically, given a voltage-response map $\Lambda$ for a $2\times 2$ rectangular electrical network, the task is to recover the non-zero, off-diagonal entries of the Kirchhoff matrix $K$. We trained a 4-layer SANN to minimize the loss $\mathcal{L}_{\text{total}}$ from equation (10). We compare this against a feed-forward neural network trained to minimize $\mathcal{L}_{\text{accuracy}}$ from equation 10) on the same data. Both networks were trained using the Adam optimizer. We also compare against a known reconstruction algorithm from Curtis and Morrow [2000] that is exact for noiseless data. The results are shown in Table 2 below. Let us shortly discuss there results: The Curtis and Morrow algorithm is an analytical method that solves the problem using theory of planar graphs and therefore, it is rigorously provable. As an analytical method, the Curtis and Morrow algorithm works also outside training data used in the other algorithms, for example with the exceptional resistor networks that have very small or large resistors, but the algorithm is very sensitive for noise. The algorithm based on a feed forward network is used in our tests as a black-box algorithm and it works well with noisy data. The SANN is between these two algorithms – it is explainable in the sense that SANN can represent the rational matrix functions, and it works moderately with noisy data. We note that in the numerical tests the training of SANNs was not fine tuned and by improving the training methods, it is possible that the results could improved.
@@
Table 2: Results of training neural networks from data to solve the network EIT problem on $2\times 2$ rectangular electrical networks. The last row (C&M) shows the reconstruction algorithm of Curtis and Morrow from Curtis and Morrow [2000].
@@ F.2 Continuous EIT problem (PDEs)
The inverse problem for electrical networks described above is a discrete version of the medical imaging problem encountered in Electrical Impedance Tomography Borcea [2002] inverse problems for finite element models for the (anisotropic) conductivity equation, and resistor networks. If $\Omega\subset\mathbb{R}^{2}$ is the interior of a simple polygon, and if one is given a triangulation of $\overline{\Omega}$, then the Finite Element Method (FEM) model for the Dirichlet problem for an elliptic partial differential equation,
@@
where $u=(u_{1}\cdots u_{N})^{t}$ corresponds to values of the FEM solution at the vertices of the triangulation, $f=(f_{1}\cdots f_{B})^{t}$ corresponds to the boundary value $f$ at the boundary vertices, and the matrix $(L^{\prime}\ L^{\prime\prime})$ depends of $\sigma$ and the triangulation and has rows whose elements add up to zero.
@@
Thus, formally, the FEM model for an elliptic partial differential equation and resistor networks lead to the same kind of equation, see Lassas et al. [2015] on details.
The Electrical Impedance Tomography (EIT) is a method proposed for medical and industrial imaging, where the objective is to determine the electrical conductivity or impedance in a medium by making voltage to current measurements on its boundary Borcea [2002]. In medical imaging it has led to new imaging modalities, for example, in monitoring lung function, see the link Pulmovista500, device by Draeger AG & Co. Recently, EIT is used in monitoring the lungs of the COVID-19 patients, see Cappellini et al. [2024].
The boundary measurements are modelled by the Dirichlet-to-Neumann map $\Lambda_{a}$, which maps the boundary value $f$ of the electric voltage the the to the electric current through the boundary (i.e., the conormal derivative) $\Lambda_{a}(f)={\nu}\cdot\sigma u|_{\partial\Omega}$.
In FEM, in the case when $\Omega$ is the interior of a simple polygon, and fix a triangulation of $\overline{\Omega}$ with vertices $\{x_{1},\ldots,x_{N}\}$ of which $\{x_{1},\ldots,x_{B}\}$ lie on $\partial\Omega$. Let $A(\Omega)$ be the set of continuous functions $\overline{\Omega}\to\mathbb{R}$ which are affine on each triangle. If $v_{k}$ is the unique function in $A(\Omega)$ which is $1$ at $x_{k}$ and 00 at all the other vertices, then $A(\Omega)=\{\sum_{k=1}^{N}c_{k}v_{k}\,;\,c_{k}\in\mathbb{C}\}$.
There is a voltage-to-current map $\Lambda_{\sigma}^{\mathrm{FEM}}$ related to the FEM model that is a disctete version of the Dirichlet-to-Neumann map. To define this map, let us consider the space $A(\partial\Omega)=A(\Omega)|_{\partial\Omega}=\{\sum_{j=1}^{B}c_{j}v_{j}|_{\partial\Omega};c_{j}\in\mathbb{R}\}$, and a quadratic form $Q_{\sigma}$ on $A(\partial\Omega)$ that is given for $f\in A(\partial\Omega)$ by
@@
where the minimum, is taken over $u\in A(\Omega)$ satisfying the boundary condition $u|_{\partial\Omega}=f$. The voltage-to-current map $\Lambda_{\sigma}^{\mathrm{FEM}}$ is a symmetric matrix $\Lambda_{\sigma}^{\mathrm{FEM}}:\mathbb{R}^{B}\to\mathbb{R}^{B}$ determined by the formula
@@
The solution of the matrix equation 91 can be using Cramer’s formula as a meromorphic function. This implies that there are a SANN that solves the FEM approximation problem equation 91 for the partial differential equation 90 and another SANN that represents the voltage-to-current map.
@@ G.1 Semialgebraic optimization
Optimization problem 92 includes as special cases polynomial optimization and quadratic/linear/mixed-integer programming. It is at least NP-hard; there is no general polynomial-time algorithm to compute $f$ from a description of $g$ and $\mathcal{S}$ [Lasserre, 2015].
Many interesting problems can be formulated in terms of 92. Below, we give a new potential application using SANNs as a hypernetwork to sparsify ReLU networks.
@@ G.2 SANNs as hypernetworks
To connect this general class of optimization problems (92) to hypernetworks, suppose we have a fixed target architecture $\mathcal{N}_{\text{target}}$ with trained weights $\theta\in\mathbb{R}^{m}$. Let $\mathcal{S}(\theta)\subset\mathbb{R}^{m}$ be the set of all weights generating equivalent networks Petersen et al. [2021]; i.e. every choice of weights from $\mathcal{S}(\theta)$ produces the same input-output pairs as $\theta$ when used in architecture $\mathcal{N}_{\text{target}}$. When $\mathcal{N}_{\text{target}}$ is a feed-forward ReLU network, $\mathcal{S}(\theta)$ is a semialgebraic set [Valluri and Campbell, 2023].
Now, different choices of objective function $g$ optimize various properties of the target network. For example, when $g$ counts the number of non-zero elements of $\theta^{*}$ (the “$L^{0}$ norm”), $f$ returns the sparsest-possible equivalent network. Using sparse $\theta^{*}$ is both more memory-efficient Bölcskei et al. [2018] and interpretable Fan et al. [2021] than using $\theta$. Likewise, $f$ can return a “nearly-equivalent” network with small Lipschitz constant by relaxing $\mathcal{S}$ and using $g$ that computes the Lipschitz constant of $\mathcal{N}_{\text{target}}$ with weights $\theta^{*}$. Reducing the Lipschitz constant of the target network improves its robustness, interpretability, and generalization performance (see Ducotterd et al. [2024] and references therein).

### v2

@@ E.3 Proof of Theorem LABEL:thm-continuous-semialgebraic-functions-computable-by-homotopy
Let $F:\mathbb{R}^{m}\to U_{F}$ be a given continuous semialgebraic function. For every ${x\in\mathbb{R}^{m}}$ there exists $H_{x,a}$ with the form (10) such that for almost every ${a\in U_{F}}$, the following hold:
@@
[…]
@@ E.4 Proof of Theorem 12
All that remains it to prove this procedure can be repeated until time $s=1$ is reached. Due to the clamping in line 5 of Algorithm 1, $z$ remains in $[-c_{\max},c_{\max}]^{n+1}$ for time $s\in[0,1]$. Since $\mathcal{N}\in ISDnet(m,n,1)$ is a piecewise polynomial, it is Lipschitz on the compact domain $\{x\}\times[-c_{\max},c_{\max}]$, which means there exists $\epsilon>0$ such that for each step $\tau$ of the $s$-glueing procedure above, we can choose $s_{\tau+1}$ such that $s_{\tau}+\epsilon<s_{\tau+1}$, and the conclusion follows. ∎
@@ E.5 Discontinuous functions
Furthermore, we carefully construct each piece of the computation to lie in disjoint $(t,s)$-space, so they can be glued together in a way consistent with a single continuous ISD network $\mathcal{N}$. This $\mathcal{N}$ computes $F$ while solving the ODE (6)–(7). Ensuring that the required $\mathcal{N}$ is indeed an ISD network (rather than an arbitrary continuous one) is one of the primary challenges in these constructions.
@@ E.6 Characteristic functions
Hypothesis 66 guarantees the sets
@@
$\texttt{ODESolve}(\mathcal{N}^{\prime}_{2},c,x,(0,1,1/2),1)=(0,1)-(0,\chi_{S_{2}}(x)).$ (80)
@@
$\texttt{ODESolve}(\mathcal{N}^{\prime}_{2},c,x,(0,0,1/2),1)=(0,0),$ (81)
@@
We can now handle the case of an arbitrary semialgebraic set $S$. The key is to exploit the “cylindrical decomposition” of $S$: every semialgebraic set in $\mathbb{R}^{m}$ is the union of finitely many disjoint sets homeomorphic to $(-1,1)^{d}$, $d\leq m$ (Bochnak et al., 1998). In particular, it is the union of finitely many disjoint locally closed sets. Since locally closed sets are the intersection of an open and a closed set, we have already shown how to represent their characteristic functions with SANNs. Furthermore, since these locally compact sets are disjoint, the characteristic function of their union can be represented by summing the individual characteristic functions with $s$-glueing. ∎
@@ E.7 Proof of Theorem 13
Every semialgebraic function has finitely many connected components (Bochnak et al., 1998), so let $\{S_{j}\}_{j=1}^{K}$ be a semialgebraic decomposition of $\mathbb{R}^{m}$ such that $F$ is continuous on each $S_{i}$.
@@
Using Lemma 42, modify each $\mathcal{N}_{i}$, $c_{i}$, so that integration occurs in the interval $s\in[1/2,1]$. Furthermore, using Lemma 43, modify each $\mathcal{N}_{i}$, $c_{i}$ to begin integration at $(0,3j).$ Finally, $t$-glue these modified $\mathcal{N}_{j}$, and $s$-glue the result to $\mathcal{N}_{0}$ at $s=1/2$. ∎
@@ E.9 SANNs always compute semialgebraic functions
In our considerations, we use the Tarski–Seidenberg theorem, which has several equivalent formulations. We use the following:
@@ Appendix F Numerical example: Inverse problem for electrical resistor networks
Suppose we are given an electrical network $\Gamma=(V,E)$ with vertices $V$ and edges $E\subset V\times V.$ Physically, an electrical network is a system of connected resistors, and it is modeled by a graph where every edge is a resistor. An inverse problem for an electrical network is to find one or all graphs and resistors on the edges when the outcome of all possible voltage and current measurements in some of vertices of the graph (called the boundary nodes) are given, see Figure 5.
We partition $V$ into boundary nodes $V_{B}$ and interior nodes $V_{I}$. Let $\gamma:E\to\mathbb{R}_{\geq 0}$ be the edge conductivities. The network response matrix (i.e. Dirichlet-to-Neumann map) $\Lambda\in\mathbb{R}^{|V_{B}|\times|V_{B}|}$ computes the current at the boundary nodes resulting from applying voltages on $V_{B}$. Our task is to recover the conductivites $\gamma$ from measurements $\Lambda$ of voltage-response on the boundary. See Curtis & Morrow (2000) for a thorough treatment of this inverse problem, and Forcey (2023) for recent developments.
@@
$K$ is a symmetric matrix with conductivities on off-diagonals and that satisfies Kirchhoff’s Law—the sum of entries in each row is 00. Curtis & Morrow (2000) showed that $\Lambda$ is the Schur complement of a particular submatrix of the Kirchhoff matrix. Precisely, we write
@@
The inverse problem is to recover an element of the preimage $F^{-1}(\{\Lambda\})$. Since preimages of semialgebraic sets (in this case the single point $\{\Lambda\}$) of semialgebraic functions ($F$) are themselves semialgebraic (Bochnak et al., 1998), we conclude $F^{-1}(\{\Lambda\})$ is a semialgebraic set.
In certain special network topologies, such as for rectangular networks (Figure 5), the preimage of $F^{-1}(\{\Lambda\})$ is a singleton, that is, the values $\gamma(i,j),$ $(i,j)\in E$ of the resistors on the graph $(V,E)$ are uniquely determined by the voltage-to-current map $\Lambda$ (Curtis & Morrow, 2000).
@@
This inverse problem has had a wide impact in several fields. In medical imaging it has led to new imaging modalities, for example, in monitoring lung function. Recently, EIT is used in monitoring the lungs of the COVID-19 patients, see Cappellini et al. (2024).
@@
We demonstrate the feasibility of training SANNs to perform non-trivial tasks using the network EIT problem as an example. Specifically, given a voltage-response map $\Lambda$ for a $2\times 2$ rectangular electrical network, the task is to recover the non-zero, off-diagonal entries of the Kirchhoff matrix $K$. We trained a 4-layer SANN to minimize the loss $\mathcal{L}_{\text{total}}$ from equation (9). We compare this against a feed-forward neural network trained to minimize $\mathcal{L}_{\text{accuracy}}$ from equation (9) on the same data. Both networks were trained using the Adam optimizer. We also compare against a known reconstruction algorithm from Curtis & Morrow (2000) that is exact for noiseless data. The results are shown in Table 2 below. Let us shortly discuss these results: The Curtis and Morrow (C&M) algorithm is an analytical method that solves the problem using the theory of planar graphs, and therefore it is provably correct for all inputs, including inputs outside the training data used for the other algorithms, such as exceptional resistor networks with very small or large resistors. However, the algorithm is very sensitive to noise. The algorithm based on a feed forward network is used in our tests as a black-box algorithm and it works well with noisy data. The SANN is between these two algorithms—it is explainable in the sense that SANN can represent the rational matrix functions, and it works moderately well with noisy data. We note that in the numerical tests the training of SANNs was not fine tuned and by improving the training methods, it is possible that the results could be improved.
@@
Table 2: Results of training neural networks from data to solve the network EIT problem on $2\times 2$ rectangular electrical networks. The last row (C&M) shows the reconstruction algorithm from Curtis & Morrow (2000).
@@ F.2 Continuous EIT problem (PDEs)
The inverse problem for electrical networks described above is a discrete version of the medical imaging problem encountered in Electrical Impedance Tomography (Borcea, 2002), an inverse problem for finite element models for the (anisotropic) conductivity equation. If $\Omega\subset\mathbb{R}^{2}$ is the interior of a simple polygon, and if one is given a triangulation of $\overline{\Omega}$, then the Finite Element Method (FEM) model for the Dirichlet problem for an elliptic partial differential equation,
@@
where $u=(u_{1}\cdots u_{N})^{t}$ corresponds to values of the FEM solution at the vertices of the triangulation, $f=(f_{1}\cdots f_{B})^{t}$ corresponds to the boundary value $f$ at the boundary vertices, and the matrix $(L^{\prime}\ L^{\prime\prime})$ depends on $\sigma$ and the triangulation and has rows whose elements add up to zero.
@@
Thus, formally, the FEM model for an elliptic partial differential equation and resistor networks lead to the same kind of equation, see Lassas et al. (2015) for details.
Electrical Impedance Tomography (EIT) is a method proposed for medical and industrial imaging, where the objective is to determine the electrical conductivity or impedance in a medium by making voltage to current measurements on its boundary (Borcea, 2002). It has led to new imaging modalities, for example the Pulmovista500 device by Draeger AG & Co. used to monitor lung function. Recently, EIT is being used to monitor the lungs of COVID-19 patients (Cappellini et al., 2024).
The boundary measurements are modeled by the Dirichlet-to-Neumann map $\Lambda_{a}$, which maps the boundary value $f$ of the electric voltage to the electric current through the boundary (i.e., the conormal derivative) $\Lambda_{a}(f)={\nu}\cdot\sigma u|_{\partial\Omega}$.
In FEM, let $\Omega$ be the interior of a simple polygon, and fix a triangulation of $\overline{\Omega}$ with vertices $\{x_{1},\ldots,x_{N}\}$ of which $\{x_{1},\ldots,x_{B}\}$ lie on $\partial\Omega$. Let $A(\Omega)$ be the set of continuous functions $\overline{\Omega}\to\mathbb{R}$ which are affine on each triangle. If $v_{k}$ is the unique function in $A(\Omega)$ which is $1$ at $x_{k}$ and 00 at all the other vertices, then $A(\Omega)=\{\sum_{k=1}^{N}c_{k}v_{k}\,;\,c_{k}\in\mathbb{C}\}$.
There is a voltage-to-current map $\Lambda_{\sigma}^{\mathrm{FEM}}$ related to the FEM model that is a discrete version of the Dirichlet-to-Neumann map. To define this map, let us consider the space $A(\partial\Omega)=A(\Omega)|_{\partial\Omega}=\{\sum_{j=1}^{B}c_{j}v_{j}|_{\partial\Omega};c_{j}\in\mathbb{R}\}$, and a quadratic form $Q_{\sigma}$ on $A(\partial\Omega)$ that is given for $f\in A(\partial\Omega)$ by
@@
where the minimum is taken over $u\in A(\Omega)$ satisfying the boundary condition $u|_{\partial\Omega}=f$. The voltage-to-current map $\Lambda_{\sigma}^{\mathrm{FEM}}$ is a symmetric matrix $\Lambda_{\sigma}^{\mathrm{FEM}}:\mathbb{R}^{B}\to\mathbb{R}^{B}$ determined by the formula
@@
The solution of the matrix equation (90) can be found using Cramer’s formula as a meromorphic function. This implies that there is a SANN that solves the FEM approximation problem (90) for the partial differential equation (89) and another SANN that represents the voltage-to-current map.
@@ G.1 Semialgebraic optimization
Optimization problem (91) includes as special cases polynomial optimization and quadratic/linear/mixed-integer programming. It is at least NP-hard; there is no general polynomial-time algorithm to compute $f$ from a description of $g$ and $\mathcal{S}$ (Lasserre, 2015).
Many interesting problems can be formulated in terms of (91). Below, we give a new potential application using SANNs as a hypernetwork to sparsify ReLU networks.
@@ G.2 SANNs as hypernetworks
To connect the general class of optimization problems (91) to hypernetworks, suppose we have a fixed target architecture $\mathcal{N}_{\text{target}}$ with trained weights $\theta\in\mathbb{R}^{m}$. Let $\mathcal{S}(\theta)\subset\mathbb{R}^{m}$ be the set of all weights generating equivalent networks (Petersen et al., 2021); i.e. every choice of weights from $\mathcal{S}(\theta)$ produces the same input-output pairs as $\theta$ when used in architecture $\mathcal{N}_{\text{target}}$. When $\mathcal{N}_{\text{target}}$ is a feed-forward ReLU network, $\mathcal{S}(\theta)$ is a semialgebraic set (Valluri & Campbell, 2023).
Now, different choices of objective function $g$ optimize various properties of the target network. For example, when $g$ counts the number of non-zero elements of $\theta^{*}$ (the “$L^{0}$ norm”), $f$ returns the sparsest-possible equivalent network. Using sparse $\theta^{*}$ is both more memory-efficient (Bölcskei et al., 2018) and interpretable (Fan et al., 2021) than using $\theta$. Likewise, $f$ can return a “nearly-equivalent” network with small Lipschitz constant by relaxing $\mathcal{S}$ and using $g$ that computes the Lipschitz constant of $\mathcal{N}_{\text{target}}$ with weights $\theta^{*}$. Reducing the Lipschitz constant of the target network improves its robustness, interpretability, and generalization performance (see Ducotterd et al. (2024) and references therein).